OpenStudy (anonymous):

How many pounds of an $83 per pound substance must be mixed with an $11 per pound substance to get 240 lbs of an alloy which will cost $28 per pound? Round to the nearest pound. Answer 57 lbs 183 lbs 205 lbs 35 lbs

OpenStudy (anonymous):

let x be the mass of $83 substance and y be the mass of $11 subtance x + y = 240 83x + 11y = 28*240

OpenStudy (anonymous):

can you solve these? x + y = 240 83x + 11y = 6720 or want to know how i get the equations?

OpenStudy (anonymous):

You Just Lost Me

OpenStudy (anonymous):

let x be the mass of $83 substance and y be the mass of $11 subtance got this? both of them have to add up to 240 pounds of the new subtance. x + y = 240 the price for 240 pounds of the new substance has to be $28 times 240 pounds. so 83$ each pounds times some pounds of the first substance + and $11 each pounds times some pounds of the second substance. 83x + 11y = 28 * 240 = 6720